OpenStudy (anonymous):

Find the mean, median, mode, and range of this data: 49, 49, 54, 55, 52, 49, 55. If necessary, round to the nearest tenth. A. 51.9, 52, 49, 6 B. 52.9, 52, 49, 104 C. 51.9, 53, 49, 6 D. 52, 52.9, 49, 104

OpenStudy (anonymous):

A. 51.9, 52, 49, 6

OpenStudy (anonymous):

To find mean\[=\frac{ 49+49+49+52+55+55+54 }{ 7 }\]\[=\frac{ 363 }{ 7 }\]\[=51.9(nearest~tenth)\]

OpenStudy (anonymous):

To find median Step 1: list the values of the data set in numerical order\[49,49,49,52,54,55,55\]Step 2 :Identify which value appears in the middle of the list.\[49,49,49,\color{red}{52},54,55,55\]\[Therefore,the~median~is~52.\]

OpenStudy (anonymous):

To find mode Step 1:Identify which value in the data set occurs most often.\[\color{red}{49},\color{red}{49},\color{red}{49},52,54,55,55\]\[Therfore,the~mode~is~49.\]

OpenStudy (anonymous):

To find range Step 1 :Biggest value-smallest value\[\color{red}{49},49,49,52,54,55,\color{red}{55}\]\[=55-49\]\[=6\]\[Therefore,the~mode~is~6.\]

OpenStudy (anonymous):

the range*

OpenStudy (anonymous):

The answer is A...correct! @Chlorophyll